### v1

# The Primacy of Magnitude in Low-Rank Adaptation

Abstract
Low-Rank Adaptation (LoRA) offers a parameter-efficient paradigm for tuning large models. While recent spectral initialization methods improve convergence and performance over the naive “Noise & Zeros” scheme, their extra computational and storage overhead undermines efficiency. In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM [^1] Code will be released at this page, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies. Our key contributions are threefold: (i) Magnitude of weight updates determines convergence. We prove low-rank structures intrinsically bound update magnitudes, unifying hyperparameter tuning in learning rate, scaling factor, and initialization as mechanisms to optimize magnitude regulation. (ii) Spectral initialization succeeds via magnitude amplification. We demystify that the presumed knowledge-driven benefit of the spectral component essentially arises from the boost in the weight update magnitude. (iii) A novel and compact initialization strategy, LoRAM, scales deterministic orthogonal bases using pretrained weight magnitudes to simulate spectral gains. Extensive experiments show that LoRAM serves as a strong baseline, retaining the full efficiency of LoRA while matching or outperforming spectral initialization across benchmarks.
1 Introduction
The rise of large pretrained models Brown2020LanguageMA ; OpenAI2023GPT4TR ; touvron2023llama ; bai2023qwen ; guo2025deepseek has driven urgent needs for parameter-efficient fine-tuning (PEFT) methods peft ; Hu2021LoRALA ; Lester2021ThePO ; He2021TowardsAU ; Edalati2022KronAPE ; zhang2025parameter . Among these, Low-Rank Adaptation (LoRA) Hu2021LoRALA stands out for its efficiency, flexibility, and stability. By freezing pretrained weights and injecting trainable low-rank matrices, LoRA enables the update of less than 1% of the parameters, significantly reducing memory and compute costs. Its plug-and-play nature achieves easy integration into diverse models, facilitating model sharing and federated learning SLoRA ; FeDeRA . Furthermore, LoRA helps prevent catastrophic forgetting biderman2024lora , making it suitable for continuous learning. These advantages have led to its wide adoption in NLP liu2022few ; Ding2022DeltaTA ; sun2023comparative ; zhao2024lora and multimodal applications ye2023mplug ; guo2023animatediff ; blattmann2023stable ; ruiz2023dreambooth ; zhang2023towards .
Despite achieved efficiency, the low-rank reparameterization constrains practical performance and convergence biderman2024lora . Besides the well-known “representation bottleneck” hyeon2021fedpara ; Edalati2022KronAPE ; Zi2023DeltaLoRAFH ; gao2024parameter ; jiang2024mora ; lialin2023relora ; xia2024chain ; PeriodicLoRA , LoRA is highly sensitive to hyperparameters due to the non-convex and non-smooth loss landscape li2024crucial . Effective training relies on careful tuning of rank zeng2023expressive ; Zhang2023AdaptiveBA ; ding2023sparse , scaling factor kalajdzievski2023rank , learning rate lora+ , initialization strategies zhu2024asymmetry ; hayou2024impact ; li2024crucial , and preconditioning li2024crucial ; zhang2024riemannian ; lora-one . Recent works increasingly leverage information from pretrained weights PiSSA ; MiLoRA ; olora or task-specific data corda ; lora-ga ; lora-one to improve the “Noise & Zeros” baseline. Among these, PiSSA PiSSA pioneers the use of Singular Value Decomposition (SVD) for LoRA initialization, employing spectral components of pretrained weights to significantly enhance convergence and performance. Subsequent works MiLoRA ; lora-ga ; olora ; lora-one ; corda extend to diverse matrix decompositions, fostering a wave of knowledge-driven initialization schemes.
While spectral initialization PiSSA ; MiLoRA ; lora-ga ; lora-one ; corda showcases considerable promise in convergence and performance, two fundamental challenges persist. First, they introduce complexity by requiring additional matrix decomposition and storage overhead, undermining usage efficiency in resource-constrained settings vera ; gao2024parameter ; QLoRA and hindering seamless integration with deep learning libraries peft ; wolf-etal-2020-transformers . Second, their success remains poorly understood. The common justification PiSSA ; MiLoRA ; corda ; mao2025survey that spectral components preserve features better than random alternatives lacks theoretical grounding. Although recent works lora-ga ; lora-one suggest that some specific initialization may approximate full-parameter gradients, the non-convex nature of LoRA renders training dynamics unpredictable.
In this paper, we demystify the knowledge-driven intuition behind spectral initialization and reveal that its effectiveness primarily stems from the magnitude scaling of weight updates. We design a minimal “Basis & Basis” initialization strategy, which demonstrates comparable performance without the overhead of SVD operations. Specifically, our key contributions can be summarized as:
  • We identify weight update magnitude as a fundamental principle for analyzing and improving LoRA’s training dynamics. This principle unifies previously independent factors, such as learning rate, scaling, and initialization, revealing their shared ability to control weight update strength and achieve comparable amplification effects when properly configured.
  • We demonstrate that initialization scheme critically shapes LoRA’s weight magnitude dynamics. Theoretically, we prove that LoRA naturally produces smaller updates than full fine-tuning, which limits its convergence and expressiveness. Moreover, we show that spectral initialization amplifies updates, providing a principled explanation for its effectiveness beyond knowledge-driven intuition.
  • Guided by the magnitude principle, we propose Magnitude-driven Initialization (LoRAM) to make LoRA initialization efficient again. LoRAM employs a logarithmic magnitude factor to retain the benefits of spectral scaling, while directly scaling deterministic orthogonal bases to eliminate the need for decomposition and storage. Extensive experiments on language and vision-language tasks establish LoRAM as a strong and practical baseline, surpassing prior initialization schemes.
Figure 1: We propose LoRAM, a magnitude-driven initialization method that enhances both the convergence and performance of LoRA while maintaining its efficiency. Unlike spectral initialization, which precomputes and stores singular components ($U,V,S$) PiSSA , LoRAM uses deterministic orthogonal bases and derives scaling from pretrained weight statistics. This elegant simplification is grounded in our analysis of LoRA through a novel lens of magnitude dynamics, where we show that the benefits of spectral values in scaling weight update magnitude can be effectively approximated.
2 Magnitude Principle for Characterizing LoRA Dynamics
2.1 Preliminaries and Notations
Given a pretrained weight matrix $W\in\mathbb{R}^{n\times m}$, LoRA Hu2021LoRALA reparameterizes the forward pass as
where $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ are trainable low-rank matrices with $r\ll\min(n,m)$, and $\alpha$ scales the update magnitude. The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for computation convenience, as illustrated in Figure 1(b). Given a loss function $L$, LoRA updates are computed as:
$\nabla_{A}L=\alpha B^{\top}\frac{\partial L}{\partial y}x^{\top}=\alpha B^{\top}(\nabla_{W}L),\quad\nabla_{B}L=\alpha\frac{\partial L}{\partial y}x^{\top}A^{\top}=\alpha(\nabla_{W}L)A^{\top}.$
Magnitude metric. To elucidate how LoRA affects the training process, we analyze the dynamics of the parameters $A$, $B$, and the resulting weight update $W_{\text{LoRA}}$. Specifically, we define the weight magnitude as $\nu[W_{\text{LoRA}}]=\frac{1}{mn}\|W_{\text{LoRA}}\|_{F}^{2}$, which serves as a central metric in our study. Assuming independent and zero-mean entries he2015delving , the expected weight magnitude is given by $\mathbb{E}[\nu[BA]]=r\,\mathbb{E}[\nu[B]]\,\mathbb{E}[\nu[A]]$. In the asymptotic regime where $m$ and $n$ are large, $\nu[W_{\text{LoRA}}]$ is approximated with the variance of $W_{\text{LoRA}}$, and $\nu[BA]\approx r\,\nu[B]\,\nu[A]$. Our analysis is motivated by the fact that LoRA introduces no change to the pretrained weights initially, i.e., $\nu[W^{(0)}_{\text{LoRA}}]=0$, while its effect emerges gradually through training. Therefore, we use the term “magnitude” instead of “variance” to highlight the cumulative growth of $W_{\text{LoRA}}$. In Appendix B, we also present a theoretical insight showing that parameter magnitude is a key determinant of LoRA’s expressiveness.
2.2 Effect of Hyperparameters on Update Magnitude
Let $\Delta W^{(t)}_{\mathrm{LoRA}}$ denote the weight update at step $t$ for LoRA framework, and $W^{(t)}_{\text{LoRA}}=\sum_{i=0}^{t-1}\Delta W^{(i)}_{\mathrm{LoRA}}$ represent cumulative adaptation. Given a learning rate $\eta$, we expand $\Delta W^{(t)}_{\mathrm{LoRA}}$ using gradient update rules, leading to the following formulations for the update magnitude [^1] See Appendix for derivation and proof:
$\Delta W^{(t)}_{\mathrm{LoRA}}$ $=\alpha\eta\left(B^{(t)}\nabla_{A}L^{(t)}+\nabla_{B}L^{(t)}A^{(t)}+\eta\nabla_{B}L^{(t)}\nabla_{A}L^{(t)}\right),$ (3)
$\text{and}\quad\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]$ $\approx r\alpha^{2}\eta^{2}\left(\nu[B^{(t)}]\nu[\nabla_{A}L^{(t)}]+\nu[\nabla_{B}L^{(t)}]\nu[A^{(t)}]\right).$ (4)
These equations indicate the complex interplay of multiple hyperparameters, distinct from the full-parameter updates given by $\Delta W^{(t)}=-\eta\nabla_{W}L^{(t)}$. We investigate the interplay among the learning rate $\eta$, scaling factor $\alpha$, and initialization magnitude, revealing a quantifiable equivalence relationship.
Proposition 1 (Parameter Scaling Equivalence).
For LoRA layers defined in Eq. (1), consider decomposing the scaling factor $\alpha=\alpha^{\prime}\alpha_{A}\alpha_{B}$, where $\alpha^{\prime},\alpha_{A},\alpha_{B}\in\mathbb{R}^{+}$. Under the commonly used optimization frameworks with negligible numerical errors, the following parametrization schemes exhibit dynamical equivalence throughout training: For all iterations $t\geq 0$, $\Delta W^{(t)}_{\mathrm{LoRA}}=\Delta\tilde{W}^{(t)}_{\mathrm{LoRA}}$ and $W^{(t)}_{\mathrm{LoRA}}=\tilde{W}^{(t)}_{\mathrm{LoRA}}$, where $\tilde{A}^{(t)}$, $\tilde{B}^{(t)}$ and $\tilde{W}^{(t)}$ represent the re-parameterized versions.
Original SGD Adam
Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
Remarks. This equivalence underscores how hyperparameters collectively regulate update magnitude, effectively reducing the search space for optimal configurations. A striking implication is that, increasing $\eta_{B}$ in LoRA+ lora+ is identical to scaling $\alpha$ in RsLoRA kalajdzievski2023rank under the "Noise & Zeros" initialization, highlighting the critical role of initialization magnitude in shaping LoRA’s training dynamics, which is rarely discussed in prior works. For the non-zero initialization, it is advisable to first adjust the initialization magnitudes and learning rates for moderate improvements, as modifying $\alpha$ inherently combines the effects of both and may result in drastic and unpredictable changes.
To demonstrate the joint effect of hyperparameters on LoRA dynamics, we conduct a controlled experiment using a 5-layer MLP with “Noise & Zeros” initialized LoRA layers, setting the intermediate dimension to 400 and the LoRA rank to 25. The network is trained on synthetic data under various hyperparameter settings, consequently using SGD and Adam optimizers with $\eta=5\times 10^{-5}$. As shown in Figure 2(a), all settings with $\alpha=16$ result in identical loss trajectories and parameter evolution, confirming the theoretical predictions. In contrast, weight updates deviate significantly when using $\alpha=1$ with $\eta$ scaled by 4, indicating that equivalence holds only under specific rules.
Figure 2: (a) Validation of Proposition 1. Each curve represents a model with unique hyperparameters. The norm difference (right axis) aggregates Frobenius norm discrepancies between the baseline model (black) and others across layers. Purple and other curves share identical learning rates but diverge due to differing initialization magnitudes. Equivalent optimization trajectories emerge from diverse hyperparameter combinations under both SGD and Adam optimizers. (b) Validation of Proposition 2. The black curve represents random orthogonal initialization. Parameter magnitudes are predominantly governed by initialization scaling, resulting in smaller weight changes compared to conventional linear layers. This necessitates the magnitude scaling in enhancing LoRA performance.
2.3 Magnitude Limitation Rooted in Low-Rank Structure
Guided by the established equivalence framework, we fix $\alpha=1$ to eliminate the interference of scaling factors, which is the most commonly-used configuration in practical implementation. In the following, we prove initialization magnitudes and other factors critically influence the weight update.
Proposition 2 (Parameter Magnitude Dynamics).
Consider LoRA parameters updated with the same learning rate $\eta$. Assume: $A^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{A}^{2}I)$, $B^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{B}^{2}I)$, $\nabla_{W}L^{(t)}\sim\mathcal{N}(\mathbf{0},\sigma_{L}^{2}I)$, and $\mathbb{E}[\langle A^{(t)},\nabla_{A}L^{(t)}\rangle]=\mathbb{E}[\langle B^{(t)},\nabla_{B}L^{(t)}\rangle]=0$. Under these conditions, the parameter magnitudes $\boldsymbol{\nu}_{t}=\left[\mathbb{E}[\nu[A^{(t)}]],\mathbb{E}[\nu[B^{(t)}]]\right]^{T}$ evolve as a linear dynamical system. Its exponential solution admits the linearized approximation under small-$\eta$ regime:
$\boldsymbol{\nu}_{t}=\left(I+\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{t-1}\approx\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix},$
where $\gamma_{A}=m\eta^{2}\sigma_{L}^{2}$, $\gamma_{B}=n\eta^{2}\sigma_{L}^{2}$. This further yields the evolution of weight update magnitude:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
Remarks.
This analysis uncovers essential properties of LoRA initialization and optimization trajectory. First, since $\gamma_{A}$ and $\gamma_{B}$ are very small values ($\ll 1$), the magnitudes of parameters $A^{(t)}$ and $B^{(t)}$ remains nearly unchanged throughout training, potentially constraining the representation capacity of the learned model. Moreover, unlike full-parameter tuning, which evolves at a linear rate of $\gamma$, the low-rank structure introduces a proportional factor $k_{1}$, significantly slowing updates. For instance, the naive “Noise & Zeros” initialization yields $k_{1}=\frac{r}{m}$, while the dimension $m$ in large models like LLaMA touvron2023llama is in the thousands or more. Despite the quadratic term accelerates growth, small gradients may temper this effect in the later training stages.
Figure 2(b) visualizes LoRA magnitude dynamics during training. We use the same network as in Figure 2(a) but apply a nonzero initialization with $\sigma_{A}=\sigma_{B}=\frac{1}{20}$. We explore a regular MLP with the same magnitude denoted as “Linear”, and a group of networks with larger initialization magnitudes. Notably, while the loss decreases significantly, the magnitudes of $A$ and $B$ remain nearly unchanged throughout training. The magnitude evolution of $W$ reveals that the basic LoRA with theoretically $k_{1}=\frac{1}{16}$, grows substantially slower than the regular MLP. Increasing the initialization magnitude effectively accelerates the growth of $W$, aligning with our theoretical analysis.
Integrating analyses in this section, we derive a magnitude principle for LoRA development:
A valid improvement to LoRA convergence will enhance weight update magnitude $\nu[W_{\text{LoRA}}]$.
As shown in Proposition 2, magnitude principle could unify and explain improvement factors in existing works, including the learning rate, scaling factor, gradients, rank and initialization schemes.
3 Demystifying Spectral Gains with Magnitude Principle
The sheer scale of modern neural networks complicates the determination of optimal LoRA initialization across layers. Drawing inspiration from recent spectral initialization methods PiSSA ; MiLoRA ; corda ; lora-ga ; lora-one , which have demonstrated improved convergence and task performance, we reinterpret their effectiveness through the lens of the magnitude principle and introduce a magnitude-driven initialization method called LoRAM. Notably, these methods requires extra SVD computations and storage, leading to increased resource overhead and implementation complexity. In contrast, LoRAM mitigates these drawbacks and achieves even better performance. In the following, we take the seminal work PiSSA PiSSA as a representative baseline and ablate other methods in experiments (see Section 4.3).
3.1 Magnitude Gain in Spectral Initialization
The PiSSA method PiSSA initializes LoRA using the spectral decomposition of pretrained weight matrix $W=USV^{\top}$, which has a rank of $\mathcal{R}[W]$. The spectral initialization is defined as:
$A^{(0)}=A_{\mathrm{SVD}}=\sqrt{S_{r}}V_{:,:r}^{\top},\quad B^{(0)}=B_{\mathrm{SVD}}=U_{:,:r}\sqrt{S_{r}},$
where $S_{r}\in\mathbb{R}^{r\times r}$ contains the top-$r$ singular values, and $U\in\mathbb{R}^{n\times n}$, $V\in\mathbb{R}^{m\times m}$ are the left and right singular vector matrices. While prior works PiSSA ; MiLoRA ; corda ; mao2025survey attribute PiSSA’s success to its ability to preserve principal components, we show that its key advantage lies in singular value weighting. By redistributing dominant variance into the initialization, PiSSA facilitates adaptive magnitude updates across layers, accelerating convergence.
Consider the statistics of the top-$r$ singular values, we define the spectral concentration factor as:
$\rho[r]\triangleq\frac{\mathbb{E}_{r}[s]^{2}}{\mathbb{E}_{\mathcal{R}[W]}[s^{2}]}=\frac{\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}}{\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}}.$
This factor captures the concentration of energy in the top-$r$ singular values. We then reformulate
$\nu(A_{\mathrm{SVD}})=\mathbb{E}_{r}[s]\nu[V_{:,:r}]=\sqrt{\frac{n\rho[r]\nu[W]}{m\mathcal{R}[W]}},\ \nu(B_{\mathrm{SVD}})=\mathbb{E}_{r}[s]\nu[U_{:,:r}]=\sqrt{\frac{m\rho[r]\nu[W]}{n\mathcal{R}[W]}}.$
Essentially, $\rho[r]$ acts as a scaling factor that redistributes variance from the pretrained weight matrix, influencing the magnitude of updates during training. Since $\rho[r]$ monotonically decreases with $r$, its impact is more pronounced for smaller $r$, making it particularly relevant for LoRA applications.
Figure 3: Illustration of spectral gain factor $Q[r]$ defined in Eq. (10) and spectral concentration factor $\rho[r]$ defined in Eq. (8) across DeBERTa-v3-base he2021debertav3 , LLaMA-2-7B touvron2023llama and FLUX.1-12B flux2024 . Values are computed from uniformly sampled layers. The white dotted line represents the linear growth rate of naive LoRA weight magnitudes, while spectral initialization exhibits faster growth. Due to its concave nature, we approximate the spectral gain factor using a logarithmic function.
Spectral Gain Factor. Taking the above magnitudes into the dynamics in Eq. (6) further derives:
$k_{1}=Q[r](m+n)\nu[W],\quad 0\leq Q[r]\triangleq\frac{\rho[r]r}{\mathcal{R}[W]}\leq 1.$
Given that $\nu[W]\sim\mathcal{O}(\min(\frac{1}{m},\frac{1}{n}))$, this results in a gain factor of at least $Q[r]$, which we term as the “spectral gain factor”. As shown in Figure 3, the spectral gain factor $Q[r]$ exhibits bounded variation in $[0,1]$ with characteristic concavity, which can be formally derived via Jensen’s inequality. Although SVD components are not completely independent preventing the theoretical monotonic increase in $Q[r]$, this concavity also suggests that the gain effect is more pronounced when $r$ is small, reinforcing the effectiveness of spectral initialization for LoRA in the parameter-efficient manner.
3.2 Efficient Magnitude-driven Initialization with LoRAM
We propose LoRAM to achieve similar magnitude update rate in Eq. (10) like PiSSA while eliminating spectral computation. As depicted in Algorithm 1, LoRAM initializes the parameter matrices as:
$A^{(0)}=A_{\mathrm{LoRAM}}=\beta\cdot\Phi_{m}^{\top},\quad B^{(0)}=B_{\mathrm{LoRAM}}=\beta\cdot\Phi_{n},\quad\beta=\left(\frac{Q[r]\cdot\nu[W]}{\nu\left[\Phi_{n}\Phi_{m}^{\top}\right]}\right)^{\frac{1}{4}}.$
Here $\Phi_{n}$ and $\Phi_{m}$ denote the first $r$ columns of an $n$- and $m$-dimensional orthogonal basis matrices, respectively. Given that $\nu[\Phi_{n}]=\frac{1}{n}$ and $\nu[\Phi_{m}]=\frac{1}{m}$, LoRAM achieves the similar magnitude as PiSSA with $k_{1}=Q[r](m+n)\nu[W].$ We can also derive $\nu[B^{(0)}A^{(0)}]=Q[r]\nu[W]$, implying LoRAM inherently ensures numerical stability and moderate corrections to the pretrained weight.
Logarithmic Gain Factor. Due to the concave nature of $Q[r]$, we approximate its analytical form using an asymptotic expansion: $Q[r]\approx\log_{\min(n,m)}(r).$ As illustrated in Figure 3, this logarithmic function effectively captures the monotonic increase nature, providing a predictable improvement than LoRA and PiSSA particularly when using a small rank $r$.
Deterministic Basis Matrix. To eliminate the need for storing initialization buffers, we adopt an analytic approach instead of random initialization. Specifically, we employ the Discrete Sine Transform (DST) basis due to its simplistic mathematical definition:
$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 1\leq i,j\leq m.$
This formulation constructs orthogonal matrices of arbitrary dimensions, ensuring reproducibility across different devices while providing provable statistical properties: $\mathbb{E}[\Phi_{m}]=0$ and $\nu[\Phi_{m}]=\frac{1}{m}$. One may wonder if randomness is required for initialization, we find it unnecessary in LoRA. In fact, DST even slightly outperforms random strategies in our ablation experiments (see Section 4.3).
Efficiency and Compatibility. Since $\beta$ and $\Phi$ avoid complex matrix operations, LoRAM retains the efficiency and storage footprint of naive LoRA. As it only modifies initialization, LoRAM remains plug-and-play, integrating seamlessly into any pipeline that supports standard LoRA. This is especially valuable for modern large models built on fixed and highly optimized frameworks. In contrast, other initialization methods require costly preprocessing, such as matrix generation lora-ga ; corda or decomposition PiSSA ; olora ; MiLoRA , which complicates adoption in standard workflows.
Input: Pretrained weight $W\in\mathbb{R}^{n\times m}$, target rank $r$
Output: Initialized parameters $A^{(0)},B^{(0)},W$
$\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r)
  // Generate basis matrices, Eq (12)
$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$
  // Compute magnitude gain factor
$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$
  // Initialize parameters
Algorithm 1 LoRAM Initialization Procedure
4 Experiments
We conduct comprehensive experiments to evaluate LoRAM efficiently implemented via the PEFT library peft . Following conventional settings PiSSA ; MiLoRA ; lora+ , we assess performance on language tasks and extend the evaluation to vision-language tasks, demonstrating LoRAM’s generalization across diverse models and modalities. All experiments are run on servers with 8 NVIDIA H800 GPUs.
Baselines. While extensive research on LoRA has explored aspects like structural modifications and rank control, these directions are largely orthogonal to our focus on hyperparameter analysis within the naive LoRA framework. In line with this, we compare LoRAM with the naive LoRA (ICLR 2022) Hu2021LoRALA , as well as several representative hyperparameter tuning strategies. We first consider weight-driven initialization (marked “§$\S$”), including PiSSA (NeurIPS 2024) PiSSA , which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA , which utilizes the last $r$ singular vectors and values; and OLoRA olora , which applies orthogonal initialization via QR decomposition. All these methods adopt a fixed scaling factor $\alpha=1$. We then include RsLoRA kalajdzievski2023rank  (marked $\dagger$”), which enhances performance by setting $\alpha=\sqrt{r}$, and LoRA+ (ICML 2024) lora+  (marked $\ddagger$”), which increases the learning rate with the recommended $\eta_{B}=4\eta_{A}$. We also evaluate data-driven initialization in the ablation study, including LoRA-GA (NeurIPS 2024) wang2024lora and CorDA (NeurIPS 2024) corda , which require extra pipeline to leverage training data information. Most of these baselines have been integrated and validated in the PEFT library.
Table 1: Comparison of LoRAM versus hyperparameter tuning baselines on NLG tasks. Experiments conducted with LLaMA2-7B model using two ranks, reporting mean $\pm$ std results (%) over three runs. Bold and underlined values represent the best and second-best performances, respectively.
Rank ##\##Param Method GSM8K MATH HumanEval MBPP Commonsense
N/A 6738M Full FT 60.34 $\pm$ 1.32 11.74 $\pm$ 0.63 32.30 $\pm$ 1.26 39.27 $\pm$ 1.01 79.20 $\pm$ 1.20
16 40M LoRA 31.51 $\pm$ 0.31 4.16 $\pm$ 0.27 15.98 $\pm$ 0.20 28.65 $\pm$ 0.47 66.56 $\pm$ 1.21
RsLoRA 39.04 $\pm$ 0.53 4.94 $\pm$ 0.40 18.85 $\pm$ 0.66 28.10 $\pm$ 0.64 73.24 $\pm$ 0.84
LoRA+ 31.69 $\pm$ 0.64 3.98 $\pm$ 0.38 18.54 $\pm$ 0.52 28.00 $\pm$ 0.81 72.19 $\pm$ 1.43
MiLoRA§ 29.70 $\pm$ 0.42 4.18 $\pm$ 0.21 14.69 $\pm$ 0.66 27.23 $\pm$ 0.53 67.90 $\pm$ 1.20
OLoRA§ 35.83 $\pm$ 0.58 4.80 $\pm$ 0.53 16.58 $\pm$ 0.38 27.44 $\pm$ 0.76 73.48 $\pm$ 1.09
PiSSA§ 37.68 $\pm$ 0.45 5.16 $\pm$ 0.41 18.37 $\pm$ 0.49 28.62 $\pm$ 0.68 73.72 $\pm$ 1.05
LoRAM§ 40.32 $\pm$ 0.43 5.30 $\pm$ 0.37 18.92 $\pm$ 0.55 28.83 $\pm$ 0.63 75.19 $\pm$ 1.10
128 320M LoRA 40.27 $\pm$ 0.70 4.72 $\pm$ 0.43 20.11 $\pm$ 0.32 28.84 $\pm$ 0.37 73.64 $\pm$ 1.13
RsLoRA 50.38 $\pm$ 0.37 7.32 $\pm$ 0.28 21.32 $\pm$ 0.70 30.73 $\pm$ 0.43 77.01 $\pm$ 1.17
LoRA+ 40.41 $\pm$ 0.67 5.28 $\pm$ 0.53 20.71 $\pm$ 0.88 29.13 $\pm$ 0.78 78.19 $\pm$ 1.33
MiLoRA§ 39.81 $\pm$ 0.89 5.18 $\pm$ 0.58 20.39 $\pm$ 0.21 29.95 $\pm$ 1.05 74.29 $\pm$ 1.09
OLoRA§ 50.10 $\pm$ 0.64 7.01 $\pm$ 0.56 20.72 $\pm$ 0.67 30.21 $\pm$ 0.89 78.61 $\pm$ 0.97
PiSSA§ 51.48 $\pm$ 0.77 7.04 $\pm$ 0.54 21.62 $\pm$ 0.48 31.07 $\pm$ 0.68 77.28 $\pm$ 0.98
LoRAM§ 51.12 $\pm$ 0.73 7.25 $\pm$ 0.68 22.03 $\pm$ 0.56 31.53 $\pm$ 0.72 77.81 $\pm$ 0.96
Table 2: Comparison of LoRAM versus hyperparameter tuning baselines on GLUE benchmark. Experiments conducted with the DeBERTa-v3-base model using rank 8, reporting mean results over three runs. Bold and underlined values represent the best and second-best performances, respectively.
Method ##\##Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
Full FT 184M 88.31 93.57 89.46 67.26 92.80 91.52 83.75 86.87
LoRA 1.33M 90.23 95.87 84.06 63.56 93.88 90.55 50.18 87.20
RsLoRA 1.33M 90.33 95.64 86.38 64.85 93.97 90.26 60.31 88.37
LoRA+ 1.33M 90.37 95.32 87.54 64.79 94.32 90.93 65.37 89.20
MiLoRA§ 1.33M 90.28 95.75 87.00 62.58 93.97 90.83 54.87 87.74
OLoRA§ 1.33M 90.19 94.83 88.72 65.59 93.36 90.74 74.09 88.52
PiSSA§ 1.33M 90.38 95.64 89.21 65.06 93.84 91.35 74.36 88.90
LoRAM§ 1.33M 90.3490.34{90.34}90.34 95.29 89.95 65.53 94.08 91.70 74.72 89.93
Table 3: Comparing LoRAM with other LoRA variants on LLaVA for multimodel tasks. Bold and underlined values indicate the top and second-best performances, respectively.
Method $\textbf{MME}_{\text{Cog}}$ $\textbf{MME}_{\text{Per}}$ MMMU AI2D ChartQA OCRBench TextVQA ScienceQA
Full FT 280 1541 0.355 0.583 0.251 0.361 0.597 0.722
LoRA 278 1402 0.331 0.557 0.231 0.333 0.536 0.684
RsLoRA 274 1385 0.334 0.573 0.227 0.328 0.539 0.694
LoRA+ 283 1389 0.341 0.565 0.229 0.331 0.545 0.690
MiLoRA§$\S$ 285 1354 0.340 0.564 0.220 0.335 0.536 0.681
OLoRA§$\S$ 288 1404 0.345 0.565 0.228 0.330 0.540 0.677
PiSSA§$\S$ 311 1411 0.344 0.564 0.232 0.338 0.547 0.686
LoRAM§$\S$ $\underline{308}$ 1406 0.350 0.571 0.238 0.336 0.551 0.700
Figure 4: Comparison of LoRA, PiSSA, and LoRAM on image customization task. Experiments conducted with the state-of-the-art FLUX.1-12B model using rank 8.
4.1 Evaluating the Performance on Natural Language Tasks
Nature Language Generation (NLG). As shown in Table 1, we conduct supervised fine-tuning of LLaMA 2-7B touvron2023llama on math, coding, and commonsense reasoning tasks. Our setup strictly follows PiSSA PiSSA , using the AdamW optimizer with a batch size of 128, a learning rate of $2\times 10^{-5}$, a warmup ratio of 0.03, and no weight decay. All experiments are performed on subsets containing 100K data points for one epoch to minimize training overhead. For math tasks, the model is tuned on MetaMathQA yu2023metamath and evaluated on GSM8K cobbe2021gsm8k and MATH yu2023metamath validation sets. For coding tasks, we use CodeFeedback zheng2024opencodeinterpreter as training dataset, with evaluations on HumanEval chen2021evaluating and MBPP austin2021program . For commonsense tasks, model is tuned on Commonsense170K hu2023llmadapters , and we report averaged accuracy on eight sub-datasets. The results in Table 1 demonstrate that LoRAM consistently outperforms LoRA variants across diverse tasks and rank settings, without requiring matrix decomposition.
Nature Language Understanding (NLU). We evaluate the NLU performance by fine-tuning the DeBERTa-v3-base model he2021debertav3 with a rank of 8 on eight tasks in the GLUE benchmark Wang2018GLUEAM . We utilize scripts from the Transformers Library wolf-etal-2020-transformers to ensure a fair comparison. All methods are trained with a learning rate of $1\times 10^{-4}$ for 3 training epochs, except for MRPC, which uses 5 epochs due to its smaller size. We report overall matched and mismatched accuracy on MNLI, Matthew’s correlation on CoLA, Pearson correlation on STS-B, and accuracy on the other datasets. As shown in Table 2, LoRAM achieves competitive performance against PiSSA across most tasks.
4.2 Evaluating the Performance on Vision-Language Tasks
Text-to-image synthesis. We adapt the advanced FLUX.1-12B flux2024 to address the image customization task, implementing LoRA, PiSSA, and LoRAM under identical configurations: a learning rate of $1\times 10^{-4}$, a batch size of 1 and 1,000 iterations. We set the rank to 8, optimizing 9.3 million parameters while maintaining computational efficiency on a single GPU. The training data and prompt template adhere to DreamBooth’s protocol ruiz2023dreambooth . Qualitative results in Figure 4 demonstrate that LoRAM exhibits marginally superior performance in detail fidelity compared to PiSSA.
Image-to-text generation. Following the pipeline of LLaVA LLaVA , we employ CLIP-ViT-L/14 radford2021learning as the visual encoder, Vicuna-13B zheng2023judging as the text decoder, and a new visual resampler guo2024llava as the connector. In the pre-training stage, we fine-tune only the perceiver resampler using the CC-595K dataset LLaVA for one epoch. During the subsequent instruction-tuning stage, we fine-tune both Vicuna and the resampler using a 656K mixture dataset LLaVA . The learning rate is set to $2\times 10^{-5}$, and the batch size is 128. We follow the official implementation and use a rank of $64$. As shown in Table 3, LoRAM achieves favorable performance across multiple multimodal benchmarks.
Training curves. We provide representative training loss curves of diverse initialization methods in Figure 5. It can be noticed that LoRAM is able to have a faster convergence rate in the early stages compared to other LoRA variants and incur smaller losses in the end.
Figure 5: Illustration of training loss curves. LoRAM achieves comparative convergence dynamics to PiSSA across diverse models and benchmarks. See tables and texts for the evaluation results.
Table 4: Results of ablation study. “Orth.” denotes the random orthogonal initialization. The left and right sides of the slash indicate the results of the method and the tracking mode, respectively. The average value is calculated over all the ranks and tasks to compare the overall trend of change.
Rank Task $\mathbf{Q[r]}$ Basis Weight-driven Data-driven + RsLoRA $\log\frac{r}{2}$ $\log r^{*}$ $\log 2r$ Orth. Gaussian MiLoRA PiSSA CorDA LoRA-GA LoRA-GA LoRAM 16 MATH 5.08 5.30 5.18 4.74 4.62 4.18 / 4.05 5.16 / 5.10 4.60 / 4.14 5.73 / 3.76 7.94 7.22 GSM8k 40.1 40.3 40.7 36.3 35.8 29.7 / 29.6 37.6 / 36.7 36.2 / 30.7 45.7 / 30.9 51.5 52.1 MBPP 28.8 28.8 28.3 28.6 27.5 27.2 / 27.8 28.6 / 29.1 25.7 / 28.3 28.3 / 27.8 33.9 31.5 HumanEval 17.1 18.9 17.1 17.7 17.7 14.6 / 14.7 18.3 / 17.3 15.2 / 15.2 19.5 / 17.7 22.0 18.3 128 MATH 7.52 7.25 7.51 7.40 7.62 5.18 / 5.04 7.04 / 6.95 6.26 / 5.04 9.18 / 7.32 9.08 11.1 GSM8k 50.7 51.1 50.4 50.2 49.8 39.8 / 39.2 51.4 / 49.5 44.5 / 40.3 54.4 / 50.8 53.6 59.4 MBPP 31.5 31.5 32.3 31.3 32.8 29.9 / 30.2 31.0 / 29.6 29.6 / 29.9 32.0 / 30.2 31.7 38.1 HumanEval 20.7 22.0 22.6 23.2 22.6 20.3 / 19.5 21.6 / 20.3 20.1 / 18.9 24.4 / 20.7 26.8 31.7 Average value 25.1 25.5 25.6 24.9 24.8 21.1 / 21.2 25.0 / 24.3 22.7 / 21.6 27.4 / 23.6 29.5 31.1
4.3 Ablating the Magnitude Principle in LoRA Improvements
We conduct ablation experiments on LLaMA-2-7B touvron2023llama under the NLG setting, focusing on the effect of magnitude gain factor, the choice of basis matrix, and the validation of the magnitude principle.
Magnitude Gain Factor. As shown in Table 4, we first evaluate different values of $Q[r]$ and observe that increasing its value slightly improves performance. We further introduce a “tracking mode” that adjusts $\beta$ in Algorithm 1 based on the initialization magnitudes from reference methods. Specifically, we set $\beta=\sqrt{\frac{\mathbb{V}[B_{\text{ref}}A_{\text{ref}}]}{\mathbb{V}\left[\Phi_{n}\Phi_{m}^{\top}\right]}}$, where $B_{\text{ref}}$ and $A_{\text{ref}}$ are initialization matrix using weight-driven PiSSA ; MiLoRA or data-driven approaches corda ; lora-ga . Under this mode, LoRAM essentially matches the performance of prior methods, confirming that magnitudes govern its performance. We also observe that PiSSA, which selects the top-$r$ singular values, outperforms MiLoRA, which selects the last $r$ singular values. This indicates that leveraging the large dominant singular values enhances performance.
Basis Matrix. We find that the choice of basis matrix generally has limited impact. For instance, replacing the DST basis with a random orthogonal or Gaussian matrix just results in only minor performance degradation. In tracking mode, substituting the SVD-derived basis with DST does not significantly impact performance. A notable exception is LoRA-GA lora-ga , which approximates the full-parameter gradient at initialization. Nonetheless, we emphasize that tracking mode fails not due to incorrectness of magnitude principle, but because the LoRA-GA matrix form maximizes gradient magnitudes of Eq. (2), making it irreplaceable by alternatives and validating magnitude principle [^2] See Proposition 5 in Appendix. We prove LoRA-GA initialization maximizes LoRA gradient magnitude..
Upper Bound of Magnitude Scaling. While increasing update magnitude generally improves performance, the benefit is not unlimited. For example, applying RsLoRA to LoRA-GA yields a clear gain at rank 8, but the improvement diminishes and may even reverse at rank 128. This suggests that magnitude scaling should be applied conservatively at higher ranks, since larger ranks inherently amplify updates, as demonstrated in our Proposition 2. Given that data-driven methods involve costly gradient and SVD computations, we recommend LoRAM with RsLoRA as a more efficient and scalable alternative for accelerating LoRA convergence and performance.
5 Conclusion
In this paper, we explore the magnitude principle of Low-Rank Adaptation (LoRA) and introduce a novel magnitude-driven initialization strategy, LoRAM, that bridges the gap between efficiency and performance. Our work demystifies the prevailing awareness surrounding spectral initialization methods, demonstrating that their success primarily stems from the amplification of weight update magnitudes. By focusing on magnitude regulation as the key driver of convergence, we provide a unified perspective that connects seemingly disparate hyperparameter adjustments, such as learning rate, scaling factor, and initialization schemes, under a single framework.
Limitations and Future Work. Despite the advancements introduced in this work, several challenges remain open for future research. First, LoRAM mimics spectral initialization magnitudes rather than seeking optimal ones; exploring alternative strategies could yield further gains. Additionally, different layers may benefit from tailored magnitude settings, motivating joint optimization with learning rate and rank. Finally, our work does not explicitly address optimization dynamics and convergence properties. These directions remain valuable for advancing parameter-efficient fine-tuning.
Appendix A Related Work
Advances in Low-Rank Adaptation.
LoRA (Hu2021LoRALA, ) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low, ; mao2025survey, ). Existing improvements to LoRA primarily focus on three directions:
Knowledge-Driven Low-Rank Initialization.
Proper initialization critically influences neural network training outcomes glorot2010understanding ; he2015delving . The standard LoRA (Hu2021LoRALA, ) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence. Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA, ; MiLoRA, ; olora, ; SLoRA, ) or task-specific data (lora-one, ; corda, ; lora-ga, ). Most prominent approaches involve Singular Value Decomposition (SVD), which allows flexible control over matrix rank. As a pioneer, PiSSA PiSSA initializes LoRA weights using principal singular components of pre-trained matrices, aligning adaptation directions with the most significant parameter variations to accelerate convergence. Subsequent works corda ; lora-ga ; lora-one ; paischer2024one propose data-driven strategies that incorporate domain knowledge into adapter construction. For instance, LoRA-GA (lora-ga, ) aligns low-rank gradient directions with full fine-tuning counterparts during initialization. While these methods boost performance and retain the core LoRA structure, their theoretical effects on optimization dynamics remain unclear. More importantly, they are less pluggable than LoRA, requiring extra computational pipelines and storage for SVD buffers.
Optimization Dynamics of LoRA. LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial, ). Most existing theoretical studies are limited to simplified and idealized scenarios, such as the lazy-training regime jang2024lora ; malladi2023kernel and infinite-width limitlora+ ; hayou2024impact . A common goal across existing studies involves ensuring feature learning stability to prevent unstable or collapsed training dynamics. For instance,  kalajdzievski2023rank proves that improper scaling factors induce gradient collapse in high-rank adapters, proposing modified scaling mechanisms to stabilize forward and backward propagation. Recent work lora+ ; zhu2024asymmetry ; hayou2024impact reveals critical asymmetries in LoRA: The two low-rank matrices exhibit divergent distinct impacts on optimization trajectories. This asymmetry motivates lora+ to employ distinct learning rates for each matrix. For SVD-based initialization, lora-one investigates principal components derived from single-step full fine-tuning gradients lora-ga , providing convergence guarantees yet under restrictive assumptions. Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA, ).
The intricate nature of LoRA and its numerous advancements motivates us to seek a streamlined principle explaining its empirical success and guiding practical applications. We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank, ; biderman2024lora, ), initialization strategies (PiSSA, ; lora-ga, ), and learning rates (lora+, ), into a cohesive perspective.
Appendix B Lower Bound on Representation Error
Proposition 1 (Lower Bound on Representation Error).
Consider the function class of LoRA-parameterized linear models:
=$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
Let $R(f)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]$ denote the expected regression loss under data distribution $\mathcal{D}$, with squared error loss $\ell(f(x),y)=\|y-f(x)\|^{2}$. Define $f^{*}=\arg\min_{f}R(f)$ as the globally optimal predictor and $f^{*}_{\mathcal{H}}=\arg\min_{f\in\mathcal{H}}R(f)$ as the optimal predictor within $\mathcal{H}$. With loss of generality, we consider the optimal predictor is linear, i.e., $f^{*}(x)=W^{*}x$. Assume that the input covariance matrix is positive definite, i.e., $\Sigma_{\mathcal{D}}=\mathbb{E}_{x\sim\mathcal{D}}[xx^{\top}]\succ 0$, and the magnitudes is limited, satisfying $\alpha r\sqrt{mnM_{1}M_{2}}<\|W^{*}-W_{0}\|_{F}$. Then, the representation error is strictly positive and lower bounded:
$R(f^{*}_{\mathcal{H}})-R(f^{*})\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\left(\|W^{*}-W_{0}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}\right)^{2}>0,$
where $\lambda_{\min}(\cdot)$ denotes the smallest eigenvalue of the matrix.
Proof.
Due to the linear learnability, we have that $y(x)=f^{*}(x)=W^{*}x$. For any $W$, we have that
$R(W)$ $=\mathbb{E}_{x\sim\mathcal{D}}\left[\|(W-W^{*})x\|^{2}\right]$ (13)
$=\mathbb{E}_{x\sim\mathcal{D}}\left[(W-W^{*})x\left((W-W^{*})x\right)^{\top}\right]$
$=\operatorname{Tr}\left((W-W^{*})\Sigma_{\mathcal{D}}(W-W^{*})^{\top}\right)$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\|W-W^{*}\|_{F}^{2}.$
The last inequality comes from $\lambda_{\min}(\Sigma_{\mathcal{D}})>0$.
Due to $\nu[A]\leq M_{1},\nu[B]\leq M_{2}$, we have that
$\|A\|_{F}^{2}\leq rmM_{1},\ \|B\|_{F}^{2}\leq rnM_{2}.$
Further, for any $A,\ B\in\mathcal{H}$, we have that
$\|W_{0}+\alpha BA-W^{*}\|_{F}$ $\geq\|W_{0}-W^{*}\|_{F}-\|\alpha BA\|_{F}$ (14)
$\geq\|W_{0}-W^{*}\|_{F}-\alpha\|A\|_{F}\|B\|_{F}$
$\geq\|W_{0}-W^{*}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}$
$>0.$
Therefore, we have that
$\quad\ R(f^{*}_{\mathcal{H}})-R(f^{*})$
$=\min_{\nu[A]\leq M_{1},\nu[B]\leq M_{2}}R(W_{0}+\alpha BA)$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\min_{\nu[A]\leq M_{1},\nu[B]\leq M_{2}}\|W_{0}+\alpha BA-W^{*}\|_{F}^{2}$
$\geq\lambda_{\min}(\Sigma_{\mathcal{D}})\left(\|W_{0}-W^{*}\|_{F}-\alpha r\sqrt{mnM_{1}M_{2}}\right)^{2}>0.$
The first inequality comes from Eq. (13) and the last inequality comes from Eq. (14). ∎
Appendix C Proof of Main Theorems
C.1 Proof of Parameter Scaling Equivalence
Proposition 2 (Parameter Scaling Equivalence).
For LoRA layers defined in Eq. (1), consider decomposing the scaling factor $\alpha=\alpha^{\prime}\alpha_{A}\alpha_{B}$, where $\alpha^{\prime},\alpha_{A},\alpha_{B}\in\mathbb{R}^{+}$. Under the commonly used optimization frameworks with negligible numerical errors, the following parametrization schemes exhibit dynamical equivalence throughout training: For all iterations $t\geq 0$, $\Delta W^{(t)}_{\mathrm{LoRA}}=\Delta\tilde{W}^{(t)}_{\mathrm{LoRA}}$ and $W^{(t)}_{\mathrm{LoRA}}=\tilde{W}^{(t)}_{\mathrm{LoRA}}$, where $\tilde{A}^{(t)}$, $\tilde{B}^{(t)}$ and $\tilde{W}^{(t)}$ represent the re-parameterized versions.
Original SGD Adam
Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
Proof.
Given LoRA’s weight decomposition $W_{\mathrm{LoRA}}=\alpha BA$ where $\alpha>0$, the gradient updates follow:
$\Delta W_{\mathrm{LoRA}}^{(t)}=\alpha(B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}),$ (15)
with learning rates $\eta_{A},\eta_{B}$ for parameters $A$ and $B$, respectively. We define reparameterized parameters $\tilde{A}=\alpha_{A}A$, $\tilde{B}=\alpha_{B}B$, and $\alpha^{\prime}=\alpha/(\alpha_{A}\alpha_{B})$. The LoRA projection becomes:
$W_{\mathrm{LoRA}}=\alpha BA=\alpha^{\prime}\tilde{B}\tilde{A}.$ (16)
This transformation preserves the functional form while redistributing scaling factors.
Proof for SGD
Under initial conditions:
$\tilde{A}^{(0)}$ $=\alpha_{A}A_{\text{init}},$ (17)
$\tilde{B}^{(0)}$ $=\alpha_{B}B_{\text{init}}.$ (18)
Gradients for reparameterized parameters:
$\nabla_{\tilde{A}}L$ $=\alpha^{\prime}\tilde{B}^{\top}\nabla_{W}L,$ (19)
$\nabla_{\tilde{B}}L$ $=\alpha^{\prime}\nabla_{W}L\tilde{A}^{\top}.$ (20)
With learning rates $\eta_{\tilde{A}}=\alpha_{A}^{2}\eta_{A}$, $\eta_{\tilde{B}}=\alpha_{B}^{2}\eta_{B}$ and $t=0$:
$\tilde{A}^{(t+1)}$ $=\tilde{A}^{(t)}-\alpha_{A}^{2}\eta_{A}\nabla_{\tilde{A}}L^{(t)}$ (21)
$=\alpha_{A}A^{(t)}-\alpha_{A}^{2}\eta_{A}(\alpha^{\prime}\alpha_{B}B^{(t)})^{\top}\nabla_{W}L^{(t)}$ (22)
$=\alpha_{A}\left(A^{(t)}-\alpha\eta_{A}(B^{(t)})^{\top}\nabla_{W}L^{(t)}\right)$ (23)
$=\alpha_{A}\left(A^{(t)}-\eta_{A}\nabla_{A}L^{(t)}\right)$ (24)
$=\alpha_{A}A^{(t+1)}.$ (25)
Similarly for $\tilde{B}$:
$\tilde{B}^{(t+1)}$ $=\alpha_{B}\left(B^{(t)}-\eta_{B}\nabla_{B}L^{(t)}\right)=\alpha_{B}B^{(t+1)}.$ (26)
These equations propagate through all $t>0$ via mathematical induction. Compare weight increments:
$\Delta\tilde{W}_{\mathrm{LoRA}}^{(t)}$ $=\alpha^{\prime}[\tilde{B}^{(t+1)}\tilde{A}^{(t+1)}-\tilde{B}^{(t)}\tilde{A}^{(t)}]$ (27)
$=\alpha^{\prime}\alpha_{A}\alpha_{B}[B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}]$ (28)
$=\alpha[B^{(t+1)}A^{(t+1)}-B^{(t)}A^{(t)}]$ (29)
$=\Delta{W}_{\mathrm{LoRA}}^{(t)}.$ (30)
Proof for Adam
Adam maintains exponential moving averages ($m_{t}$, $v_{t}$) for gradients. Under parameter scaling $\tilde{P}=kP$, gradients transform as $\nabla_{\tilde{P}}L=k\nabla_{P}L$. The momentum terms inherit scaling factors:
$m_{\tilde{P}}$ $=\beta_{1}m_{\tilde{P}}+(1-\beta_{1})k\nabla_{P}L,$ (31)
$v_{\tilde{P}}$ $=\beta_{2}v_{\tilde{P}}+(1-\beta_{2})k^{2}(\nabla_{P}L)^{2}.$ (32)
The pratical gradients for Adam:
$\nabla^{\dagger}\tilde{P}$ $=m_{\tilde{P}}/\sqrt{v_{\tilde{P}}+\epsilon}$ (33)
$=[\beta_{1}m_{P}+...]/\sqrt{\beta_{2}v_{P}+...}$ (34)
When setting $\epsilon=0$ for alleviating numerical errors, we have $\nabla^{\dagger}\tilde{P}=\nabla^{\dagger}P$. Setting $\eta_{\tilde{P}}=k\eta_{P}$ cancels scaling factors, preserving update magnitudes. With learning rates $\eta_{\tilde{A}}=\alpha_{A}\eta_{A}$, $\eta_{\tilde{B}}=\alpha_{B}\eta_{B}$ and $t=0$:
$\tilde{A}^{(t+1)}$ $=\tilde{A}^{(t)}-\alpha_{A}\eta_{A}\nabla^{\dagger}_{\tilde{A}}L^{(t)}$ (35)
$=\alpha_{A}A^{(t)}-\alpha_{A}\eta_{A}\nabla^{\dagger}_{A}L^{(t)}$ (36)
$=\alpha_{A}A^{(t+1)}.$ (37)
Similarly for $\tilde{B}$. These equalities propagate through all $t>0$ via mathematical induction.
C.2 Proof of Proposition 2
For the convenience, we split the proof of Proposition 2 into two parts.
Proposition 3 (Parameter Magnitude Dynamics).
Consider LoRA parameters $A^{(t)}$ and $B^{(t)}$ updated with learning rate $\eta$. Assume: $A^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{A}^{2}I)$, $B^{(0)}\sim\mathcal{N}(\mathbf{0},\sigma_{B}^{2}I)$, $\nabla_{W}L\sim\mathcal{N}(\mathbf{0},\sigma_{L}^{2}I)$, and $\mathbb{E}[\langle A^{(t)},\nabla_{A}L^{(t)}\rangle]=\mathbb{E}[\langle B^{(t)},\nabla_{B}L^{(t)}\rangle]=0$. Under these conditions, the parameter norm vector $\boldsymbol{e}_{t}=\left[\mathbb{E}[\|A^{(t)}\|_{F}^{2}],\mathbb{E}[\|B^{(t)}\|_{F}^{2}]\right]^{T}$ evolve as the following linear dynamical system (left) with closed-form solution (right):
$\boldsymbol{e}_{t+1}=\left(I+\begin{bmatrix}0&\gamma_{A}\\ \gamma_{B}&0\end{bmatrix}\right)\boldsymbol{e}_{t},\quad\boldsymbol{e}_{t}=\begin{bmatrix}\frac{\lambda_{+}^{t}+\lambda_{-}^{t}}{2}&\sqrt{\frac{\gamma_{A}}{\gamma_{B}}}\frac{\lambda_{+}^{t}-\lambda_{-}^{t}}{2}\\ \sqrt{\frac{\gamma_{B}}{\gamma_{A}}}\frac{\lambda_{+}^{t}-\lambda_{-}^{t}}{2}&\frac{\lambda_{+}^{t}+\lambda_{-}^{t}}{2}\end{bmatrix}\boldsymbol{e}_{0}.$ (38)
where $\gamma_{A}=m\eta^{2}\sigma_{L}^{2}$, $\gamma_{B}=n\eta^{2}\sigma_{L}^{2}$ and $\lambda_{\pm}=1\pm\sqrt{\gamma_{A}\gamma_{B}}$. Similarly, the parameter magnitudes $\boldsymbol{\nu}_{t}=\left[\mathbb{E}[\nu[A^{(t)}]],\mathbb{E}[\nu[B^{(t)}]]\right]^{T}$ evolve as a linear dynamical system:
$\boldsymbol{\nu}_{t}=\left(I+\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{t-1}.$
Proof.
Following the gradient descent update rule, the parameter dynamics are:
$\begin{cases}A_{t+1}=A_{t}-\eta\nabla_{A}L_{t}\\ B_{t+1}=B_{t}-\eta\nabla_{B}L_{t}\end{cases},$ (40)
where the gradient relationships $\nabla_{A}L=B^{\top}(\nabla_{W}L)$ and $\nabla_{B}L=(\nabla_{W}L)A^{\top}$ are derived from the LoRA architecture. Expand the Frobenius norm for $A_{t+1}$:
$\|A_{t+1}\|_{F}^{2}=\|A_{t}\|_{F}^{2}-2\eta\langle A_{t},\nabla_{A}L_{t}\rangle+\eta^{2}\|\nabla_{A}L_{t}\|_{F}^{2}.$ (41)
Taking expectations (using independence $\mathbb{E}[\langle A_{t},\nabla_{A}L_{t}\rangle]=0$):
$\mathbb{E}\|A_{t+1}\|_{F}^{2}=\mathbb{E}\|A_{t}\|_{F}^{2}+\eta^{2}\mathbb{E}\|\nabla_{A}L_{t}\|_{F}^{2}.$ (42)
Similarly, for $B_{t+1}$:
$\mathbb{E}\|B_{t+1}\|_{F}^{2}=\mathbb{E}\|B_{t}\|_{F}^{2}+\eta^{2}\mathbb{E}\|\nabla_{B}L_{t}\|_{F}^{2}.$ (43)
Expand the norm of $\nabla_{A}L=B^{\top}(\nabla_{W}L)$:
$\|\nabla_{A}L\|_{F}^{2}=\text{Tr}((\nabla_{A}L)(\nabla_{A}L)^{\top})=\text{Tr}(B^{\top}(\nabla_{W}L)(\nabla_{W}L)^{\top}B).$ (44)
Taking expectations (using gradient entry independence):
$\mathbb{E}\|\nabla_{A}L\|_{F}^{2}=\text{Tr}\left(B^{\top}\mathbb{E}[(\nabla_{W}L)(\nabla_{W}L)^{\top}]B\right)=m\sigma_{L}^{2}\|B\|_{F}^{2}.$ (45)
where $\mathbb{E}[(\nabla_{W}L)(\nabla_{W}L)^{\top}]=m\sigma_{L}^{2}I_{n}$ follows from the i.i.d. assumption on gradient entries.
For $\nabla_{B}L=(\nabla_{W}L)A^{\top}$:
$\mathbb{E}\|\nabla_{B}L\|_{F}^{2}=\text{Tr}\left(A\mathbb{E}[(\nabla_{W}L)^{\top}(\nabla_{W}L)]A^{\top}\right)=n\sigma_{L}^{2}\|A\|_{F}^{2}.$ (46)
Substitute gradient norms results:
$\begin{cases}\mathbb{E}\|A_{t+1}\|_{F}^{2}=\mathbb{E}\|A_{t}\|_{F}^{2}+\eta^{2}m\sigma_{L}^{2}\mathbb{E}\|B_{t}\|_{F}^{2}\\ \mathbb{E}\|B_{t+1}\|_{F}^{2}=\mathbb{E}\|B_{t}\|_{F}^{2}+\eta^{2}n\sigma_{L}^{2}\mathbb{E}\|A_{t}\|_{F}^{2}\end{cases}.$ (47)
Define parameters and state vector:
$\gamma_{A}=\eta^{2}m\sigma_{L}^{2},\quad\gamma_{B}=\eta^{2}n\sigma_{L}^{2},\quad\boldsymbol{e}_{t}=\begin{bmatrix}\mathbb{E}\|A_{t}\|_{F}^{2}\\ \mathbb{E}\|B_{t}\|_{F}^{2}\end{bmatrix}.$ (48)
The recursive system becomes:
$\boldsymbol{e}_{t+1}=\begin{bmatrix}1&\gamma_{A}\\ \gamma_{B}&1\end{bmatrix}\boldsymbol{e}_{t}=\left(I+\begin{bmatrix}0&\gamma_{A}\\ \gamma_{B}&0\end{bmatrix}\right)\boldsymbol{e}_{t}.$ (49)
Similar results for $\boldsymbol{\nu}_{t}$ can be obtained using the same proof techniques.
Proposition 4 (Linearized Dynamics Approximation).
For sufficiently small learning rate $\eta$, the closed-form solution admits the first-order approximation:
$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0},\quad\text{where }\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix}.$ (50)
This yields the update variance expansion under small-$\eta$ regime:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+k_{2}\gamma^{2}t^{2},$ (51)
with $\gamma=\eta^{2}\sigma_{L}^{2}$, $k_{1}=r(m\nu[A^{(0)}]^{2}+n\nu[B^{(0)}]^{2})$, and $k_{2}=rmn\nu[A^{(0)}]\nu[B^{(0)}]$.
Proof.
For a sufficiently small learning rate $\eta$, the product $\gamma_{A}\gamma_{B}=mn\eta^{4}\sigma_{L}^{4}$ is very small. Hence, the eigenvalues
$\lambda_{\pm}=1\pm\sqrt{\gamma_{A}\gamma_{B}}$ (52)
can be approximated via a first-order Taylor expansion. Thus, for small $t$, the closed-form solution for $\boldsymbol{\nu}_{t}$ can be approximated as
$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0}.$ (53)
By definition, we set
$\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\,\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\,\gamma_{A}\sigma_{A}^{2}\end{bmatrix},$ (54)
which agrees with the first-order expansion of the linear system. ∎
Appendix D Analyzing LoRA-GA from Magnitude Principle
Proposition 5 (LoRA-GA initialization maximizes low-rank gradient magnitude).
Given a gradient matrix $\nabla_{W}L\in\mathbb{R}^{n\times m}$ with SVD decomposition $\nabla_{W}L=USV^{\top}$, the optimal rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$ in Eq. (2), under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, are given by:
$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r},$ (55)
where $V_{:,:r}$ and $U_{:,:r}$ contain the first $r$ right and left singular vectors of $\nabla_{W}L$, respectively.
Proof.
We seek to determine the rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$, subject to the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$. From Eq. (2), the LoRA gradient update follows:
$\nabla_{A}L=\nabla_{W}LB^{\top},\quad\nabla_{B}L=\nabla_{W}^{\top}LA^{\top}.$ (56)
By substituting the singular value decomposition of $\nabla_{W}L$, given by
$\nabla_{W}L=USV^{\top},$ (57)
we obtain
$\nabla_{A}L=USV^{\top}B^{\top},\quad\nabla_{B}L=VSU^{\top}A^{\top}.$ (58)
Taking the Frobenius norm and using the cyclic property of the trace, we have
$\|\nabla_{A}L\|_{F}^{2}=\text{tr}(BS^{2}B^{\top}),\quad\|\nabla_{B}L\|_{F}^{2}=\text{tr}(AS^{2}A^{\top}).$ (59)
To maximize these expressions under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, we use the variational characterization of the singular value decomposition. The optimal choice is to align $A$ and $B$ with the dominant singular vectors of $\nabla_{W}L$, leading to
$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r}.$ (60)
This choice ensures that the gradient magnitude is maximized while satisfying the given constraints, completing the proof. ∎
Appendix E Details on Key Formulas
E.1 Weight Update Magnitude
$\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]\approx r\alpha^{2}\eta^{2}\left(\nu[B^{(t)}]\nu[\nabla_{A}L^{(t)}]+\nu[\nabla_{B}L^{(t)}]\nu[A^{(t)}]\right).$
Proof.
Given the LoRA parameter update rule:
$\Delta W^{(t)}_{\mathrm{LoRA}}=\alpha\eta\left(B^{(t)}\nabla_{A}L^{(t)}+\nabla_{B}L^{(t)}A^{(t)}\right)+\mathcal{O}(\eta^{2}),$ (61)
where we retain first-order terms in $\eta$ under small learning rate assumption.
Assumptions:
  • (A1)
    Independence: $B\perp\!\!\!\perp\nabla_{A}L$, $\nabla_{B}L\perp\!\!\!\perp A$
  • (A2)
    Zero-mean initialization: $\mathbb{E}[B_{ij}]=\mathbb{E}[A_{kl}]=0$
  • (A3)
    Spatial homogeneity: $var[B_{ij}]=\nu[B]$, $var[(\nabla_{A}L)_{ij}]=\nu[\nabla_{A}L]$.
For entry $(m,n)$ in $\Delta W^{(t)}_{\mathrm{LoRA}}$:
$var\left[(\Delta W)_{mn}\right]$ $\approx\alpha^{2}\eta^{2}var\left[\sum_{k=1}^{r}\left(B_{mk}(\nabla_{A}L)_{kn}+(\nabla_{B}L)_{mk}A_{kn}\right)\right]$ (62)
$=\alpha^{2}\eta^{2}\left({var\left[\sum_{k}B_{mk}(\nabla_{A}L)_{kn}\right]}+var\left[\sum_{k}(\nabla_{B}L)_{mk}A_{kn}\right]\right).$ (63)
By assumptions (A1)-(A3):
$var\left[\sum_{k}B_{mk}(\nabla_{A}L)_{kn}\right]$ $=\sum_{k}var[B_{mk}]var[(\nabla_{A}L)_{kn}]$ (64)
$=r\nu[B]\nu[\nabla_{A}L].$ (65)
Similarly:
$var\left[\sum_{k}(\nabla_{B}L)_{mk}A_{kn}\right]$ $=r\nu[\nabla_{B}L]\nu[A].$ (66)
Combining these terms:
$\nu[\Delta W^{(t)}_{\mathrm{LoRA}}]$ $\approx\alpha^{2}\eta^{2}\left(r\nu[B]\nu[\nabla_{A}L]+r\nu[\nabla_{B}L]\nu[A]\right)$ (67)
$=r\alpha^{2}\eta^{2}\left(\nu[B]\nu[\nabla_{A}L]+\nu[\nabla_{B}L]\nu[A]\right).$ (68)
where $\nu[B]$, $\nu[A]$ inherit their magnitudes from initialization scheme, and $\nu[\nabla L]$ terms reflect task-specific loss landscape characteristics. ∎
E.2 Spectral Concentration Factor
Let the singular values of the pretrained weight matrix $W$ be $\{s_{i}\}_{i=1}^{\mathcal{R}[W]}$. Define the average of the top-$r$ singular values as
$\mathbb{E}_{r}[s]=\frac{1}{r}\sum_{i=1}^{r}s_{i},$ (69)
and the average of the squares of all singular values as
$\mathbb{E}_{\mathcal{R}[W]}[s^{2}]=\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}.$ (70)
Because the square function is convex, Jensen’s inequality implies
$\mathbb{E}_{r}[s]^{2}\leq\mathbb{E}_{r}[s^{2}],$ (71)
with equality if and only if all $s_{i}$ (for $i=1,\dots,r$) are equal. Moreover, since lower singular values generally contribute less to the overall energy, as $r$ increases the value of $\mathbb{E}_{r}[s]$ decreases relative to $\mathbb{E}_{\mathcal{R}[W]}[s^{2}]$, making $\rho[r]$ a monotonically decreasing function of $r$.
E.3 Variance of $A_{\mathrm{SVD}}$ and $B_{\mathrm{SVD}}$
Recall that the PiSSA initialization is given by
$A_{\mathrm{SVD}}=\sqrt{S_{r}}\,V_{:,:r}^{\top},\quad B_{\mathrm{SVD}}=U_{:,:r}\sqrt{S_{r}},$ (72)
where $S_{r}$ is a diagonal matrix containing the top-$r$ singular values of $W$. Assuming that the columns of $V$ (and similarly, $U$) form an orthonormal basis, the variance of $V_{:,:r}$ (taken elementwise) is approximately $\nu[V_{:r}]\approx\frac{1}{m}$ (or $\frac{1}{n}$ for $U$), since for an orthogonal matrix the energy is uniformly distributed. Thus, the variance of $A_{\mathrm{SVD}}$ can be expressed as:
$\nu(A_{\mathrm{SVD}})=\frac{\mathrm{Tr}(\sqrt{S^{T}_{r}}V_{:,:r}^{T}V_{:,:r}\sqrt{S})}{mr}=\frac{\sum_{i=1}^{r}s_{i}}{mr}=\mathbb{E}_{r}[s]\nu[V_{:r}].$ (73)
To connect this with the variance of $W$, note that
$\nu[W]=\frac{1}{mn}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}.$ (74)
Thus, we can relate $\mathbb{E}_{r}[s]$ to $\nu[W]$ via the spectral concentration factor $\rho[r]$. Incorporating a factor of $n$ to account for the dimensions of $A_{\mathrm{SVD}}$, we obtain:
$\nu(A_{\mathrm{SVD}})=\sqrt{\frac{n\,\rho[r]\,\nu[W]}{m\,\mathcal{R}[W]}}.$ (75)
A similar argument, with the roles of $m$ and $n$ interchanged, leads to the expression for $\nu(B_{\mathrm{SVD}})$.
E.4 Spectral Gain Factor
The dynamics of the LoRA weight update variance are captured by an expression of the form:
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+k_{2}\gamma^{2}t^{2},$ (76)
where $k_{1}$ is the linear evolution rate. Substituting the variance expressions derived for $A_{\mathrm{SVD}}$ and $B_{\mathrm{SVD}}$ into the dynamical system (see Eq. (6)), we obtain:
$k_{1}=\frac{\rho[r]\,r\,(m+n)}{\mathcal{R}[W]}\,\nu[W].$ (77)
Defining
$Q[r]\triangleq\frac{\rho[r]\,r}{\mathcal{R}[W]},$ (78)
this expression becomes:
$k_{1}=Q[r](m+n)\nu[W],$ (79)
with the constraint $0\leq Q[r]\leq 1$. When $\nu[W]\sim\mathcal{O}(\min(1/m,1/n))$, the factor $Q[r]$ effectively quantifies the amplification of the weight update magnitude due to the spectral initialization.
Proof.
Recall that
$\rho[r]\triangleq\frac{\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}}{\frac{1}{\mathcal{R}[W]}\sum_{i=1}^{\mathcal{R}[W]}s_{i}^{2}},$ (80)
and the spectral gain factor is defined as
$Q[r]\triangleq\frac{\rho[r]\,r}{\mathcal{R}[W]}.$ (81)
Let $m=\mathcal{R}[W]$. By Jensen’s inequality (or by the Cauchy–Schwarz inequality), we have
$\left(\frac{1}{r}\sum_{i=1}^{r}s_{i}\right)^{2}\leq\frac{1}{r}\sum_{i=1}^{r}s_{i}^{2}.$ (82)
Therefore,
$\rho[r]\leq\frac{\frac{1}{r}\sum_{i=1}^{r}s_{i}^{2}}{\frac{1}{m}\sum_{i=1}^{m}s_{i}^{2}}=\frac{m}{r}\cdot\frac{\sum_{i=1}^{r}s_{i}^{2}}{\sum_{i=1}^{m}s_{i}^{2}}\leq\frac{m}{r},$ (83)
since $\frac{\sum_{i=1}^{r}s_{i}^{2}}{\sum_{i=1}^{m}s_{i}^{2}}\leq 1$. Substituting this bound into the definition of $Q[r]$ yields
$Q[r]=\frac{\rho[r]\,r}{m}\leq\frac{m}{r}\cdot\frac{r}{m}=1.$ (84)
This completes the proof that $Q[r]\leq 1$. ∎

### v2

@@
 While recent spectral initialization methods improve convergence and performance over the naive “Noise & Zeros” scheme, their extra computational and storage overhead undermines efficiency.
-In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM [^1] Code will be released at this page, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies.
+In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies [^1] Code is available here.
 Our key contributions are threefold: (i) Magnitude of weight updates determines convergence.
@@
 Its plug-and-play nature achieves easy integration into diverse models, facilitating model sharing and federated learning SLoRA; FeDeRA.
-Furthermore, LoRA helps prevent catastrophic forgetting biderman2024lora , making it suitable for continuous learning.
-These advantages have led to its wide adoption in NLP liu2022few ; Ding2022DeltaTA ; sun2023comparative ; zhao2024lora and multimodal applications ye2023mplug ; guo2023animatediff ; blattmann2023stable ; ruiz2023dreambooth ; zhang2023towards .
+Additionally, LoRA helps prevent catastrophic forgetting biderman2024lora, making it well-suited for continual learning.
+These advantages have led to its wide adoption in multilingual NLP liu2022few; Ding2022DeltaTA; sun2023comparative; zhao2024lora and multimodal applications ye2023mplug; guo2023animatediff; blattmann2023stable; ruiz2023dreambooth.
 Despite achieved efficiency, the low-rank reparameterization constrains practical performance and convergence biderman2024lora.
@@
 Given a pretrained weight matrix $W\in\mathbb{R}^{n\times m}$, LoRA Hu2021LoRALA reparameterizes the forward pass as
+$y=Wx+W_{\text{LoRA}}x=Wx+\alpha(BA-B^{(0)}A^{(0)})x,$
 where $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ are trainable low-rank matrices with $r\ll\min(n,m)$, and $\alpha$ scales the update magnitude.
-The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for computation convenience, as illustrated in Figure 1(b).
+The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for the convenience, as illustrated in Figure 1(b).
 Given a loss function $L$, LoRA updates are computed as:
@@
 Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
-Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
-[…]
+Initialization $\begin{aligned} A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
+Learning Rates $\begin{aligned} \eta_{A}>0,\eta_{B}>0\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha^{2}_{A}\eta_{A},\eta_{\tilde{B}}=\alpha^{2}_{B}\eta_{B}\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha_{A}\eta_{A},\eta_{\tilde{B}}=\alpha_{B}\eta_{B}\end{aligned}$
 Remarks.
@@
 where $\gamma_{A}=m\eta^{2}\sigma_{L}^{2}$, $\gamma_{B}=n\eta^{2}\sigma_{L}^{2}$. This further yields the evolution of weight update magnitude:
-$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
+$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L}^{2},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
 Remarks.
@@
 Specifically, we employ the Discrete Sine Transform (DST) basis due to its simplistic mathematical definition:
-$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 1\leq i,j\leq m.$
+$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 0\leq i,j<m.$
 This formulation constructs orthogonal matrices of arbitrary dimensions, ensuring reproducibility across different devices while providing provable statistical properties: $\mathbb{E}[\Phi_{m}]=0$ and $\nu[\Phi_{m}]=\frac{1}{m}$.
@@
 Output: Initialized parameters $A^{(0)},B^{(0)},W$
-$\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r)
-  // Generate basis matrices, Eq (12)
-$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$
-  // Compute magnitude gain factor
-$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$
-  // Initialize parameters
+0.2cm $\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r) $\triangleright$ Generate basis matrices, e.g., Eq (12)
+$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$ $\triangleright$ Compute magnitude gain factor
+$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$ $\triangleright$ Initialize parameters
 Algorithm 1 LoRAM Initialization Procedure
@@
 In line with this, we compare LoRAM with the naive LoRA (ICLR 2022) Hu2021LoRALA, as well as several representative hyperparameter tuning strategies.
-We first consider weight-driven initialization (marked “§$\S$”), including PiSSA (NeurIPS 2024) PiSSA , which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA , which utilizes the last $r$ singular vectors and values; and OLoRA olora , which applies orthogonal initialization via QR decomposition.
+We first consider weight-driven initialization (marked $\S$”), including PiSSA (NeurIPS 2024) PiSSA, which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA, which utilizes the last $r$ singular vectors and values; and OLoRA olora, which applies orthogonal initialization via QR decomposition.
 All these methods adopt a fixed scaling factor $\alpha=1$.
@@
 Bold and underlined values represent the best and second-best performances, respectively.
-Rank ##\##Param Method GSM8K MATH HumanEval MBPP Commonsense
+Rank #\#Param Method GSM8K MATH HumanEval MBPP Commonsense
 N/A 6738M Full FT 60.34 $\pm$ 1.32 11.74 $\pm$ 0.63 32.30 $\pm$ 1.26 39.27 $\pm$ 1.01 79.20 $\pm$ 1.20
@@
 Bold and underlined values represent the best and second-best performances, respectively.
-Method ##\##Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
+Method #\#Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
 Full FT 184M 88.31 93.57 89.46 67.26 92.80 91.52 83.75 86.87
@@
 PiSSA§ 1.33M 90.38 95.64 89.21 65.06 93.84 91.35 74.36 88.90
-LoRAM§ 1.33M 90.3490.34{90.34}90.34 95.29 89.95 65.53 94.08 91.70 74.72 89.93
+LoRAM§ 1.33M 90.34{90.34} 95.29 89.95 65.53 94.08 91.70 74.72 89.93
 Table 3: Comparing LoRAM with other LoRA variants on LLaVA for multimodel tasks. Bold and underlined values indicate the top and second-best performances, respectively.
@@
 LoRA+ 283 1389 0.341 0.565 0.229 0.331 0.545 0.690
-MiLoRA§$\S$ 285 1354 0.340 0.564 0.220 0.335 0.536 0.681
-OLoRA§$\S$ 288 1404 0.345 0.565 0.228 0.330 0.540 0.677
-PiSSA§$\S$ 311 1411 0.344 0.564 0.232 0.338 0.547 0.686
-LoRAM§$\S$ $\underline{308}$ 1406 0.350 0.571 0.238 0.336 0.551 0.700
+[…]
+[…]
+[…]
+[…]
 Figure 4: Comparison of LoRA, PiSSA, and LoRAM on image customization task. Experiments conducted with the state-of-the-art FLUX.1-12B model using rank 8.
@@
 Advances in Low-Rank Adaptation.
-LoRA (Hu2021LoRALA, ) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low, ; mao2025survey, ).
+The growing scale (Brown2020LanguageMA; OpenAI2023GPT4TR; touvron2023llama; bai2023qwen; karras2020analyzing; rombach2022high) of large pre-trained models across various domains urgently demands efficient fine-tuning methods (zhao2024lora; ye2023mplug; guo2023animatediff; blattmann2023stable; ruiz2023dreambooth; goga; ei2; li2025top; han2025joyagents).
+LoRA (Hu2021LoRALA) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low; mao2025survey).
 Existing improvements to LoRA primarily focus on three directions:
+  • The low-rank constraint creates bottlenecks when learning complex features.
+Recent approaches enhance expressiveness through iterative stacking of LoRA modules (lialin2023relora; xia2024chain; Zi2023DeltaLoRAFH; PeriodicLoRA), high-order matrix operations (hyeon2021fedpara; Edalati2022KronAPE), and customized architectural designs (journals/corr/abs-2402-17263; jiang2024mora; li2024loran; zhong2024neat; Flora; Zhang2023AdaptiveBA; liu2024dora).
+  • The non-convex landscapes pose challenges for numerical optimization and hyperparameter tuning (li2024crucial).
+Researchers have explored to improve scaling factor (kalajdzievski2023rank; Sebastian2023; biderman2024lora), learning rate (lora+), dropout rate (lora-dropout), optimizers (li2024crucial; zhang2024riemannian; lora-one; lora-ga), and initialization strategies (PiSSA; MiLoRA; olora; hayou2024impact; lora-one; corda; lora-ga).
+  • Recent works reduce computational and memory overhead through freezing parameters (lora-fa; zhu2024asymmetry), designing more compact adapters (vera; lora-xs; gao2024parameter) and parameter quantization (PiSSA; QLoRA; LoftQ).
 Knowledge-Driven Low-Rank Initialization.
 Proper initialization critically influences neural network training outcomes glorot2010understanding; he2015delving.
-The standard LoRA (Hu2021LoRALA, ) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence.
-Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA, ; MiLoRA, ; olora, ; SLoRA, ) or task-specific data (lora-one, ; corda, ; lora-ga, ).
+The standard LoRA (Hu2021LoRALA) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence.
+Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA; MiLoRA; olora; SLoRA) or task-specific data (lora-one; corda; lora-ga).
 Most prominent approaches involve Singular Value Decomposition (SVD), which allows flexible control over matrix rank.
@@
 Subsequent works corda; lora-ga; lora-one; paischer2024one propose data-driven strategies that incorporate domain knowledge into adapter construction.
-For instance, LoRA-GA (lora-ga, ) aligns low-rank gradient directions with full fine-tuning counterparts during initialization.
+For instance, LoRA-GA (lora-ga) aligns low-rank gradient directions with full fine-tuning counterparts during initialization.
 While these methods boost performance and retain the core LoRA structure, their theoretical effects on optimization dynamics remain unclear.
@@
 Optimization Dynamics of LoRA.
-LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial, ).
+LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial).
 Most existing theoretical studies are limited to simplified and idealized scenarios, such as the lazy-training regime jang2024lora; malladi2023kernel and infinite-width limitlora+; hayou2024impact.
@@
 For SVD-based initialization, lora-one investigates principal components derived from single-step full fine-tuning gradients lora-ga, providing convergence guarantees yet under restrictive assumptions.
-Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA, ).
+Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA).
 The intricate nature of LoRA and its numerous advancements motivates us to seek a streamlined principle explaining its empirical success and guiding practical applications.
-We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank, ; biderman2024lora, ), initialization strategies (PiSSA, ; lora-ga, ), and learning rates (lora+, ), into a cohesive perspective.
+We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank; biderman2024lora), initialization strategies (PiSSA; lora-ga), and learning rates (lora+), into a cohesive perspective.
 Appendix B Lower Bound on Representation Error
@@
 Consider the function class of LoRA-parameterized linear models:
-=$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
+$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
 Let $R(f)=\mathbb{E}_{(x,y)\sim\mathcal{D}}[\ell(f(x),y)]$ denote the expected regression loss under data distribution $\mathcal{D}$, with squared error loss $\ell(f(x),y)=\|y-f(x)\|^{2}$.
@@
 Representation $\alpha BAx$ $\alpha^{\prime}\tilde{B}\tilde{A}x$ $\alpha^{\prime}\tilde{B}\tilde{A}x$
-Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
-[…]
+Initialization $\begin{aligned} A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
+Learning Rates $\begin{aligned} \eta_{A}>0,\eta_{B}>0\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha^{2}_{A}\eta_{A},\eta_{\tilde{B}}=\alpha^{2}_{B}\eta_{B}\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha_{A}\eta_{A},\eta_{\tilde{B}}=\alpha_{B}\eta_{B}\end{aligned}$
 Proof.
@@
 For sufficiently small learning rate $\eta$, the closed-form solution admits the first-order approximation:
-$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0},\quad\text{where }\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix}.$ (50)
+$\boldsymbol{\nu}_{t}\approx\left(I+t\begin{bmatrix}0&\gamma_{B}\\ \gamma_{A}&0\end{bmatrix}\right)\boldsymbol{\nu}_{0},\quad\text{where }\boldsymbol{\nu}_{t}\triangleq\begin{bmatrix}\nu[A^{(t)}]\\ \nu[B^{(t)}]\end{bmatrix}=\begin{bmatrix}\sigma_{A}^{2}+t\gamma_{B}\sigma_{B}^{2}\\ \sigma_{B}^{2}+t\gamma_{A}\sigma_{A}^{2}\end{bmatrix}$.
+(50)
 This yields the update variance expansion under small-$\eta$ regime:
@@
 Proof.
-We seek to determine the rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$, subject to the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$.
-From Eq.
-(2), the LoRA gradient update follows:
-$\nabla_{A}L=\nabla_{W}LB^{\top},\quad\nabla_{B}L=\nabla_{W}^{\top}LA^{\top}.$ (56)
-By substituting the singular value decomposition of $\nabla_{W}L$, given by
-$\nabla_{W}L=USV^{\top},$ (57)
-we obtain
-$\nabla_{A}L=USV^{\top}B^{\top},\quad\nabla_{B}L=VSU^{\top}A^{\top}.$ (58)
-Taking the Frobenius norm and using the cyclic property of the trace, we have
-$\|\nabla_{A}L\|_{F}^{2}=\text{tr}(BS^{2}B^{\top}),\quad\|\nabla_{B}L\|_{F}^{2}=\text{tr}(AS^{2}A^{\top}).$ (59)
-To maximize these expressions under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, we use the variational characterization of the singular value decomposition.
-The optimal choice is to align $A$ and $B$ with the dominant singular vectors of $\nabla_{W}L$, leading to
-$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r}.$ (60)
-This choice ensures that the gradient magnitude is maximized while satisfying the given constraints, completing the proof. ∎
+Let $G=\nabla_{W}L\in\mathbb{R}^{n\times m}$ and write its SVD $G=USV^{\top}$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$. From the LoRA gradient expressions we have
+$\nabla_{A}L=B^{\top}G,\qquad\nabla_{B}L=GA^{\top}.$
+Note that $\nabla_{A}L$ does not depend on $A$ and $\nabla_{B}L$ does not depend on $B$, hence the two maximization problems decouple.
+(i) Maximizing $\|\nabla_{B}L\|_{F}^{2}$ w.r.t. $A$ under $\|A\|_{F}^{2}=r$. Let $A^{\prime}:=AV$. Then $\|A^{\prime}\|_{F}^{2}=\|A\|_{F}^{2}=r$ and
+$\|GA^{\top}\|_{F}^{2}=\|US(A^{\prime})^{\top}\|_{F}^{2}=\|S(A^{\prime})^{\top}\|_{F}^{2}=\mathrm{tr}\big(A^{\prime}S^{2}(A^{\prime})^{\top}\big).$
+Form the Lagrangian $\mathcal{L}(A^{\prime},\lambda)=\mathrm{tr}(A^{\prime}S^{2}(A^{\prime})^{\top})-\lambda(\mathrm{tr}(A^{\prime}(A^{\prime})^{\top})-r)$.
+Taking derivative w.r.t.
+$A^{\prime}$ yields the stationarity condition $A^{\prime}S^{2}=\lambda A^{\prime}$.
+Thus the rows of $A^{\prime}$ lie in the eigenspaces of $S^{2}$.
+To maximize the trace, one places the Frobenius norm onto the coordinates corresponding to the largest diagonal entries of $S^{2}$, i.e.
+the first $r$ singular values.
+Taking $A^{\prime\star}=[I_{r}\mid 0]$ attains the maximum
+$\max_{\|A\|_{F}^{2}=r}\|GA^{\top}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2},$
+and $A^{\star}=A^{\prime\star}V^{\top}=V_{:,:r}^{\top}$ (up to orthonormal transformations within the chosen $r$-subspace).
+(ii) Maximizing $\|\nabla_{A}L\|_{F}^{2}$ w.r.t. $B$ under $\|B\|_{F}^{2}=r$. Analogously let $B^{\prime}:=U^{\top}B$. Then
+$\|B^{\top}G\|_{F}^{2}=\mathrm{tr}\big((B^{\prime})^{\top}S^{2}B^{\prime}\big),$
+and the same Lagrange argument gives that the optimal choice concentrates norm on the first $r$ coordinates, producing
+$\max_{\|B\|_{F}^{2}=r}\|B^{\top}G\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2},$
+and one may take $B^{\star}=U_{:,:r}$.
+Hence the stated solutions $A^{\star}=V_{:,:r}^{\top}$ and $B^{\star}=U_{:,:r}$ are optimal, with the above maximal values. This completes the proof. ∎
 Appendix E Details on Key Formulas
@@
 This completes the proof that $Q[r]\leq 1$. ∎
+Appendix F Further Clarifications
+In this section, we provide additional details regarding the experimental setup for our theoretical validations and justify the core assumptions underlying our propositions.
+F.1 Clarification on Figure 2: Experimental Setup and Optimizer Dynamics
+The experiments illustrated in Figure 2 were conducted in a controlled environment to isolate and validate our theoretical claims regarding hyperparameter equivalence and magnitude dynamics.
+The setup uses a 5-layer MLP with an intermediate dimension of 400, where LoRA modules are trained to fit a mapping for randomly generated synthetic data.
+This simple setting effectively tests the fundamental fitting capabilities of LoRA.
+Figure 2(a) is specifically designed to empirically validate Proposition 1, reproduced with clearer separation in Figure 6.
+We deliberately use two different optimizers, SGD for the first 2,500 steps and Adam thereafter, to demonstrate that when the hyperparameter product $\alpha^{\prime}\alpha_{A}\alpha_{B}$ is held constant, LoRA exhibits equivalent training trajectories regardless of the optimizer.
+This equivalence is confirmed by the overlapping loss curves.
+Furthermore, the norm difference $\Delta\tilde{W}_{\text{LoRA}}^{(t)}-\Delta W_{\text{LoRA}}^{(t)}||_{F}^{2}$ between the baseline and other equivalent settings remains near zero, confirming that the learned weights themselves evolve identically.
+Figure 6: A detailed view of the validation of Proposition 1, separating the SGD and Adam optimization phases.
+Each curve represents a model with a unique hyperparameter combination.
+The norm difference (right axis) aggregates the Frobenius norm discrepancies between the baseline model (black curve) and others across all layers.
+The results show that diverse hyperparameter sets can produce identical optimization trajectories, confirming our theoretical equivalence framework.
+F.2 Clarification on Figure 3: Direct Relationship to PiSSA
+The goal of Figure 3 is to visualize the source of magnitude gain in spectral initialization methods, for which we use PiSSA PiSSA as a representative example. The link is direct:
+  • The solid colored curves for the spectral concentration factor ($\rho[r]$) and spectral gain factor ($Q[r]$) are calculated directly from the SVD of the pretrained weights.
+This is precisely the mechanism that PiSSA employs for initialization.
+Therefore, these curves illustrate the inherent magnitude dynamics of a PiSSA-initialized model.
+  • The plots reveal two key insights from our analysis of PiSSA.
+First, the plot of $\rho[r]$ shows that spectral energy is highly concentrated in the top singular values, explaining why PiSSA is particularly effective at low ranks.
+Second, the plot of $Q[r]$ quantifies the significant magnitude gain that PiSSA (colored lines) achieves over the naive "Noise & Zeros" baseline (white dotted line).
+This is precisely the gain that LoRAM is designed to mimic efficiently without performing SVD.
+Appendix G Further Analysis on the Interplay of Initialization, Learning Rate, and Performance
+Our primary experiments were conducted under a unified hyperparameter configuration to ensure a fair and controlled comparison.
+However, different initialization strategies may achieve optimal performance under varying hyperparameters, particularly the learning rate.
+To provide a more nuanced understanding, we conducted a supplementary analysis investigating the performance of LoRAM, PiSSA PiSSA, and MiLoRA MiLoRA under different learning rates.
+Specifically, we compare the results from our original experiments using a moderate learning rate ($2\times 10^{-5}$) with new results obtained using a higher learning rate ($2\times 10^{-4}$), which aligns with the setting used in the original MiLoRA paper.
+The detailed results for ranks $r=16$ and $r=128$ are presented in Table 5 and Table 6, respectively.
+Table 5: Performance of PiSSA, LoRAM, and MiLoRA with different learning rates at rank $r=16$.
+Learning rate $2\times 10^{-5}$ $2\times 10^{-4}$
+Method PiSSA LoRAM MiLoRA PiSSA LoRAM MiLoRA
+GSM8K 37.68 40.32 29.70 52.46 53.29 46.62
+MATH 5.16 5.30 4.18 8.80 8.92 6.18
+HumanEval 18.37 18.92 14.69 24.47 25.62 17.14
+MBPP 28.62 28.83 27.23 31.22 31.74 28.65
+Commonsense 73.72 75.19 67.90 77.02 79.11 76.67
+Table 6: Performance of PiSSA, LoRAM, and MiLoRA with different learning rates at rank $r=128$.
+Learning rate $2\times 10^{-5}$ $2\times 10^{-4}$
+Method PiSSA LoRAM MiLoRA PiSSA LoRAM MiLoRA
+GSM8K 51.48 51.12 39.81 58.37 59.28 54.66
+MATH 7.04 7.25 5.18 11.46 10.76 9.20
+HumanEval 21.62 22.03 20.39 26.81 29.95 28.02
+MBPP 31.07 31.53 29.95 36.54 37.80 33.35
+Commonsense 77.28 77.81 74.29 67.09 74.23 79.01
+The results reveal several key patterns:
+  • At a lower rank ($r=16$), all methods generally benefit from the higher learning rate, showing improved performance across tasks.
+This suggests that in low-rank settings, a larger learning rate can enhance convergence speed.
+  • At a higher rank ($r=128$), a clear divergence emerges.
+While most methods still improve, PiSSA exhibits a notable performance degradation on the Commonsense dataset.
+Its training loss curve in this high-rank, high-LR configuration plateaus early, suggesting that the amplified updates overshoot the effective descent direction.
+  • These findings align perfectly with our magnitude principle.
+PiSSA’s use of principal singular components leads to stronger initial magnitude amplification, which necessitates a smaller optimal learning rate to maintain stability.
+Conversely, MiLoRA, which initializes with smaller minor singular components, can benefit from a larger learning rate.
+This analysis reinforces our claim from Section 4.3: “Magnitude scaling should be applied conservatively at higher ranks, since larger ranks inherently amplify updates.”
+This supplementary analysis underscores that the optimal learning rate is intrinsically linked to the magnitude scaling introduced by the initialization method.
+Appendix H Efficiency Analysis: Computational and Memory Overhead
+A primary motivation for LoRAM is to achieve the performance gains of spectral initialization methods while preserving the computational efficiency and minimal memory footprint of the original LoRA framework.
+In this section, we provide a detailed comparison of the theoretical and practical overhead associated with LoRAM, LoRA, and PiSSA.
+H.1 Theoretical Complexity
+The theoretical time and space complexities for the initialization phase of each method are summarized in Table 7.
+LoRA’s overhead is minimal, stemming from random matrix initialization.
+LoRAM maintains this same linear complexity, as its deterministic DST basis is generated efficiently through vector multiplication.
+In contrast, PiSSA’s complexity is dominated by the SVD of the pretrained weight matrices, a computationally intensive operation.
+Furthermore, PiSSA requires storing both the original and decomposed components, effectively doubling the space requirement compared to LoRA and LoRAM.
+Table 7: Theoretical initialization time and space complexities of LoRA, LoRAM, and PiSSA.
+Method LoRA LoRAM PiSSA
+Time $O(mr+nr)$ $O(mr+nr)$ $O(\min(m^{2}n,mn^{2}))$
+Space $O(r(m+n))$ $O(r(m+n))$ $O(2r(m+n))$
+H.2 Practical Performance
+We empirically validated these complexities by measuring the practical initialization time and memory usage while fine-tuning LLaMA-7B on an 8-GPU server.
+The results, shown in Table 8, highlight two distinct workflows for PiSSA:
+  1. 1.
+    Pre-processing : This approach first computes and saves the residual model after subtracting the low-rank approximation.
+While the initialization itself is faster, this process incurs a substantial one-time storage cost, requiring over 12.5 GB to store the residual weights in addition to the LoRA adapter weights.
+  2. 2.
+    Direct Workflow: This method computes SVD on-the-fly, which avoids the large storage overhead.
+However, it is significantly slower (often taking over 10 minutes in our setup) due to known bottlenecks related to CPU-based SVD computations before GPU transfer .
+As shown in the table, LoRAM’s practical performance is nearly identical to that of the standard LoRA implementation, demonstrating its exceptional efficiency .
+It successfully eliminates the significant time and space overhead introduced by SVD-based methods like PiSSA, confirming its utility as a lightweight yet powerful initialization strategy.
+Table 8: Practical initialization time and space cost of LoRA, LoRAM, and PiSSA on LLaMA-7B.
+Metric LoRA LoRAM PiSSA (pre-process) PiSSA (direct)
+Time ($r=16$) 1.36s 0.95s 51.32s >> 10min
+Time ($r=128$) 4.23s 2.19s 57.79s >> 10min
+Time ($r=512$) 8.50s 5.19s 110.31s >> 10min
+Space ($r=16$) 152MB 152MB 12.5GB + 152MB 305MB
+Space ($r=128$) 1.2GB 1.2GB 12.5GB + 1.2GB 2.4GB
+Space ($r=512$) 4.8GB 4.8GB 12.5GB + 4.8GB 9.6GB
+Appendix I Scope and Limitations of the Magnitude Principle
+While our work establishes the magnitude of weight updates as a fundamental driver of LoRA’s performance, it is crucial to clearly define the scope and limitations of this principle.
+Our central claim is that magnitude plays a primary, not universal, role in the success of various hyperparameter-tuning strategies.
+The principle provides a coherent and predictive lens for unifying these strategies, rather than suggesting that magnitude is the sole determinant of all outcomes.
+The applicability and effects of magnitude scaling are subject to several important considerations:
+  • Interplay with Other Hyperparameters: As indicated by our analysis in Proposition 2, the magnitude principle is not limited to initialization magnitudes ($\sigma_{A}^{2},\sigma_{B}^{2}$) alone.
+It is intrinsically linked to other key factors, including the LoRA rank ($r$), the learning rate ($\eta$), and the task-dependent gradient variance ($\sigma_{L}^{2}$).
+The final performance is a result of the complex interplay among all these components.
+  • Interaction with Rank Size: A key finding, highlighted in our ablation studies and consistent with our theoretical analysis, is that the benefits of magnitude scaling are not monotonic.
+We empirically observe that the performance gains from increased magnitude scaling tend to diminish and can even reverse at higher ranks.
+This is because larger ranks inherently amplify updates, as predicted by Proposition 2, suggesting that magnitude should be scaled more conservatively in high-rank settings.
+  • On Optimality: Our work aims to identify and validate update magnitude as a core mechanism influencing LoRA’s training dynamics, thereby demystifying the success of methods like PiSSA.
+We do not claim to have identified an optimal magnitude scaling strategy.
+Determining the optimal magnitude for different models, tasks, and layers remains a challenging and important direction for future research.
+In summary, the magnitude principle serves as a powerful analytical tool for understanding and designing LoRA-based methods, but its application should be contextualized within the broader optimization landscape.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main claims in the abstract and introduction have clearly reflected the theoretical and empirical contributions of the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We have declared the limitations in the article, including some theoretical assumptions, and have cited relevant literature to show that these assumptions are commonly used in the field.
+We have also included additional technical limitations at the end of the main paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We have provided full proofs for each proposition in the Appendix, and for the important formulas in the paper, we also give the corresponding derivation in the Appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: All benchmarks, as well as the experimental setup, are from or based on work that has been open-sourced.
+We have given relevant references in the paper and have given specific hyperparameter settings.
+Moreover, our proposed LoRAM does not require tuning parameters, as illustrated in Algorithm 1.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: Since we have conducted extensive experiments on several datasets, providing detailed scripts is intricate. The code will be made available upon acceptance of the article.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All benchmarks, as well as the experimental setup, are from or based on work that has been open-sourced.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We have provided the variance of the test results, and also the range of fluctuations in the training curves.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39.
+Justification: Our experiments are based on a server with 8 H800 GPUs.
+The timing of experiments is diverse depending on the datasets, while there will be no differences between LoRAM and other approaches on the same dataset, since they do not modify the training process.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have carefully checked and followed the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: Our paper presents a general fine-tuning technique that is not specific to actual application scenarios.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We have properly cited the original paper that produced the code package or dataset.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: The paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.